OpenStudy (abby09sj09):

sin(cos^-1(x)) rewrite the expression as an algebraic expression in x. help please! :)

OpenStudy (3mar):

Are you familiar with trigonometric function?

OpenStudy (abby09sj09):

as in SOHCAHTOA or something else?

OpenStudy (3mar):

what is SOHCAHTOA ?

OpenStudy (abby09sj09):

sin = opp/hyp, cos = adj/hyp, tan = opp/adj

OpenStudy (3mar):

seem this is for easy saving!

OpenStudy (abby09sj09):

sorry, i don't understand what you're saying

OpenStudy (3mar):

I mean it is for easy remember/restore!

OpenStudy (abby09sj09):

oh, ok. yes!

OpenStudy (3mar):

What do you think about the problem?

OpenStudy (abby09sj09):